OpenStudy (anonymous):

The legs of a right triangle measure 3x and 15. If the hypotenuse measures 3x + 3, what is the value of x?

jimthompson5910 (jim_thompson5910):

 a^2 + b^2 = c^2 (3x)^2 + 15^2 = (3x + 3)^2 9x^2 + 225 = 9x^2 + 18x + 9 225 = 18x + 9 I'll let you finish up

OpenStudy (anonymous):

x=12 (: Thanks!

jimthompson5910 (jim_thompson5910):

you got it